One mole of Nitrogen gas and 3 mole Hydrogen are kept in a container of capacity 8.21 litre at 300k to produce NH3. Find the partial pressure of Nitrogen gas and Hydrogen gas if partial pressure of NH3 after sufficient time was found to be 3atm. 

Grade:11

1 Answers

Sahani Kumar
5 years ago
Given :
Number of moles of Nitrogen(N2) = 1mole
Number of moles of Hydrogen (H2) = 3moles  
So total number of moles =1+3= 4moles 
Total pressure = 3 atm. 
Now,  
        Mole fraction of N2=( number of moles of N2)/(total no of moles) 
So, 
Mole fraction of N2=1/4=0.25
 Now, 
        Partial pressure of N2=(mole fraction of N2)*(total pressure) 
On putting values we get
Partial pressure of N2=( 0.25) * (3) =0.75 atm
Regard 
Askiitian member, 
If you like this answer please approve this answer to encourage us. Thank you.